OpenStudy (anonymous):

How do you simplify 2y/2y to make it y^2?

OpenStudy (solomonzelman):

2y/2y is not equal to y^2, it is equal to 1

OpenStudy (mokeira):

@SolomonZelman am I right? or have I forced the answer lol

OpenStudy (solomonzelman):

No

OpenStudy (solomonzelman):

thee second thing you wrote, \(\LARGE\color{black}{ \frac{1}{2}y }\) is wrong.

OpenStudy (mokeira):

i thought so too

OpenStudy (mokeira):

\[2y \times \frac{ 1 }{2y }=1\]

OpenStudy (solomonzelman):

Yes, because \(\LARGE\color{black}{ \frac{1}{2y} ≠\frac{1}{2}y }\)

OpenStudy (solomonzelman):

yes it is 1... number over itself is one.

OpenStudy (solomonzelman):

btw, \(\LARGE\color{blue}{ v^0=v^{b-b}= \frac{v^b}{v^b} =1 }\)